# A concise tutorial review of Reverse Osmosis and Electrodialysis

### v5

@@ 5.2 Flow of solutes through membranes according to solution-friction theory
To describe the flow of ions through a porous tortuous medium, we first of all set up an expression for the chemical potential of an ion. In this section the porous medium is a membrane, so these two terms are used interchangeably. We can consider multiple contributions to the chemical potential of an ion (unit J/mol) in a porous medium, adding up to [23]
@@
where the reference value $\overline{\mu}_{\text{ref},i}$ is of importance when ions take part in chemical reactions, while $RT\,\ln\left(c_{i}/c\mathrm{{}_{ref}}\right)$ relates to diffusion (ion entropy), with $c\mathrm{{}_{ref}}$ a reference concentration, for instance, $c\mathrm{{}_{ref}}=1$ mM. Furthermore, $z_{i}FV$ relates to electrostatic effects, $\overline{\mu}_{\text{aff},i}$ to a chemical affinity, leading to (one contribution to) a solubility $S_{i}$ or partitioning coefficient $\Phi_{i}$, and $\overline{\mu}_{\text{exc},i}$ describes volumetric interactions between ions, and between ions and the porous medium [22]. When only interactions of an ion with the porous medium are of importance (and not between ions), then $\overline{\mu}_{\text{exc},i}$ is a constant factor that can be combined with $\overline{\mu}_{\text{aff},i}$.
We can now set up Eq. (49) for an ion inside and outside the membrane, equate the two expressions, and for a neutral solute arrive for the distribution of a solute across the solution/membrane interface at
$\frac{c_{\text{m},i}}{c_{\infty,i}}=\exp\left(-\tfrac{1}{RT}\,\Delta\left(\overline{\mu}_{\text{aff},i}+\overline{\mu}_{\text{exc},i}\right)\right)=S_{i}=\Phi_{i}$ (50)
where the solubility $S_{i}$, or partitioning coefficient $\Phi_{i}$, is a constant factor if the chemical affinity and excess terms do not depend (too much) on concentrations in solution and in the membrane, and where $\Delta$ refers to a difference between in the membrane and in solution [16]. The excess contribution is often described as function of an ion size-pore size ratio, $\lambda$, described for a perfectly cylindrical pore and spherical ions by $\Phi_{i}=\left(1-\lambda_{i}\right)^{2}$, but many, more accurate, expressions are also available [55]. For ions we must extend Eq. (50) to account for their charge, and we then end up with the extended Boltzmann equation, Eq. (34).
@@
where we consider a single coordinate, x. An extension to multiple coordinates or dimensions is straightforward. We can leave out here a gradient-term related to $\overline{\mu}_{\text{aff},i}$ because this ion-membrane affinity will be invariant across the membrane. And for the same reason we also leave out the excess-term that relates to volume effects. Thus we only have to consider a force related to diffusion, and to the electric field, i.e., electromigration.
@@
$\mathcal{F}_{\text{friction},i}=-\sum_{j}f_{i\text{-}j}^{*}\left(v_{i}-v_{j}\right)=-f_{i\text{-w}}^{*}\left(v_{i}-v\mathrm{{}_{w}}\right)-f_{i\text{-m}}^{*}\;v_{i}$ (53)
where water velocity is $v\mathrm{{}_{w}}$ and we included in the last part that the velocity of the membrane is zero, $v\mathrm{{}_{m}}\!=\!0$. We present here velocities as superficial velocities, per unit total membrane area. A more detailed analysis is the two-fluid model (TFM) which starts at interstitial velocities, and includes pore tortuosity, and also includes how the space occupied by ions is not available for water, as explained in ref. [55].
We can write the ion-water frictional coefficient as the inverse of a diffusion coefficient of the ion in the membrane, $D_{\text{m},i}$, which is the Einstein equation, $f_{i\text{-w}}^{*}=RT/{D_{\text{m},i}}$ (p. 128 in ref. [40]). This diffusion coefficient is lower than in free solution because it includes the porosity and tortuosity of the membrane pores. We now obtain an extended Nernst-Planck equation for the molar flux of ions inside a membrane (Eq. (7) in ref. [5]) where convection, diffusion, and electromigration, all contribute [50, 35]
$J_{i}=K_{\text{f},i}c_{\text{m},i}v\mathrm{{}_{w}}-K_{\text{f},i}D_{\text{m},i}\left(\frac{\partial c_{\text{m},i}}{\partial x}+z_{i}c_{\text{m},i}\frac{\partial\phi}{\partial x}\right)$ (54)
and where we also incorporated that $J_{i}=c_{\text{m},i}v_{i}$. Concentrations here are those inside the membrane pores. The friction factor, or hindrance function, $K_{\text{f},i}$, is given by $K_{\text{f},i}=1/\left\{1+f_{i\text{-m}}^{*}/f_{i\text{-w}}^{*}\right\}$, which has a value between 0 and 1, and describes that ions not only have friction with the water that directly envelopes them, but also with the membrane matrix. Eq. (54) is a general expression for ion transport in porous media, extended compared to the standard Nernst-Planck equation with convection and with ion-membrane friction, which leads to the factor $K_{\text{f},i}$.
@@ 5.3 Flow of water through membranes according to solution-friction theory
We continue with a description of transport of water through the pores of a porous medium. As discussed before, in SF theory we assume the pores are filled with water completely, and ions are assumed to be without volume, i.e., in the theory they are what is called ‘point charges’. A full account including solute volume effects, is the two-fluid model (TFM), see ref. [55]. We can now set up a balance of forces acting on a volume element of water in the membrane pores, with the driving force a gradient in solvent pressure, i.e., $-\partial P^{\text{s}}/\partial x$ is the force acting on a volume element of water. It indeed has the correct unit N/m3. The water has friction with the membrane structure, $-f_{\text{w-m}}RT\left(v\mathrm{{}_{w}}-v\mathrm{{}_{m}}\right)$, where $v\mathrm{{}_{m}}\!=\!0$ because the membrane is immobile (ref. [40], p. 126), and water has friction with all types of solutes, and this friction is proportional to the concentration of that type of solute, which leads to $-\sum_{i}^{\text{ }}f_{i\text{-w}}^{*}c_{\text{m},i}\left(v\mathrm{{}_{w}}-v_{i}\right)$. The summation of all these forces is set to zero, resulting in [35]
$-\frac{\partial P^{\text{s}}}{\partial x}=f_{\text{w-m}}RTv\mathrm{{}_{w}}+\sum_{i}f_{i\text{-w}}^{*}c_{\text{m},i}\left(v\mathrm{{}_{w}}-v_{i}\right)\,.$ (55)
@@
We can solve Eq. (55) at each position in the membrane, together with the extended Nernst-Planck equation, Eq. (54). Now, a significant simplification is arrived at if we assume there is no ion-membrane friction, and thus we can set $K_{\text{f},i}\!=\!1$ in the NP equation, Eq. (55), for all ions. Then inserting the NP equation in Eq. (55) results in
$-\frac{1}{RT}\left(\frac{\partial P^{\text{h}}}{\partial x}-\frac{\partial\Pi}{\partial x}\right)=f_{\text{w-m}}v\mathrm{{}_{w}}+\sum_{i}\left(\frac{\partial c_{\text{m},i}}{\partial x}+z_{i}c_{\text{m},i}\frac{\partial\phi}{\partial x}\right)$ (56)
@@
where we implemented that the solvent pressure $P^{\text{s}}$ is given by $P^{\text{s}}=P^{\text{h}}-\Pi$, and $D_{\text{m},i}=RT/f_{i\text{-w}}^{*}$. Because in the ideal case that we discuss here the osmotic pressure is the total ion concentration times RT, thus $\Pi=RT\cdot\sum_{i}c_{\text{m},i}$, Eq. (56) simplifies to
$-\frac{1}{RT}\frac{\partial P^{\text{h}}}{\partial x}=f_{\text{w-m}}v\mathrm{{}_{w}}-X\,\frac{\partial\phi}{\partial x}$ (57)
@@
and thus we have an exact cancellation of the osmotic pressure acting on the water, and the diffusional forces acting on the ions, which transfer to the water via ion-water friction. But this only occurs when the ions have no friction with other phases other than the water. In the derivation of Eq. (57) we also included local electroneutrality in the membrane, $\sum_{i}z_{i}c_{\text{m},i}+X=0$. We can rewrite Eq. (57) to [30, 48]
[…]
@@
The final element of transport theory through a porous structure filled with a continuum fluid such as water, with dissolved solutes, is mechanical equilibrium at the interface between the porous medium and the outside solution. This was already discussed in section 3. Here we derive the mechanical equilibrium condition based on Eq. (56), because with all gradient terms much larger in a thin interfacial region than the terms proportional to velocities (or in other words, the structure of the interfacial region is not significantly modified by flows of ions and water across it), we obtain the result that the solvent pressure $P^{\text{s}}=P^{\text{h}}-\Pi$ is invariant across any thin interface [34, 48, 44], and this leads to the earlier conclusion that the change in osmotic pressure $\Pi$ (either up or down) equals the change in hydrostatic pressure $P^{\text{h}}$ across an interface.
@@ 5.4 Counterfluxes of salt and water in osmosis
To illustrate SF theory, and to show how well it works, we discuss results of an accurate fit of SF theory to data of the flow of water and salt across an ion-exchange membrane. This experiment is called osmosis, and water and salt flow because of a difference in salt concentration between the two outside solutions on either side of the membrane, without pressure differences or current. Results of this experiment are presented in Fig. 7 as function of the concentration on the low-salinity side, $c\mathrm{{}_{d}}$, which has values changing (in time) from 4 mM to 130 mM. During that time the salt concentration on the concentrate side decreases from 800 to 525 mM. Over the 35 hr of the experiment, the volume on the diluate side drops by 40% because water flows to the concentrated side, see ref. [55] for details. Thus, in this experiment water flows to the high-salinity side, while salt flows in the other direction. Consequently, inside the membrane these two flows are in opposite direction. This makes this an experiment that is not easy to reproduce accurately in a theoretical model, i.e., it is a very good test of a mass transport theory.
@@
When we apply SF theory, with cation and anion fluxes the same (which is because the electrical current through the membrane is zero), and with equal diffusion coefficients of the ions in the membrane, we find that it works excellently, and results in the very good fit to the data shown in Fig. 7. The salt flux, $J\mathrm{{}_{salt}}$, follows from solving Eq. (54) while implementing that the flux of cations is equal to that of anions, and then we integrate. This leads to Eq. (5) in ref. [3] which we solve assuming $c\mathrm{{}_{T,m}}$ changes linearly, after which we implement Eq. (46). Here we need a value of $\Phi_{i}$ which we base on ref. [63], from which we derive a value of $\Phi_{i}\!=\!0.82$. The water flux is calculated based on Eq. (55) which after integration results in Eq. (S17) in ref. [3]. Membrane charge density is $X\!=\!-5.1$ M. To fit the model to the data, we derive $k\mathrm{{}_{m}}\!=\!2.25$ LMH (for a membrane thickness of $L\mathrm{{}_{m}}=120$ $\mu$m, this corresponds to a diffusion coefficient in the membrane about 20$\times$ less than in free solution), $K_{\text{f},i}\!=\!0.50$, and for the water permeability we use $A\!=\!9.2$ mLMH/bar. Values for $k\mathrm{{}_{m}}$ and $A$ are in line with earlier reported values for ED membranes in ref. [51]. Thus, we can conclude that SF theory, which includes chemical and mechanical equilibrium on solution/membrane interfaces, is very well capable of describing fluxes of water and salt in the osmosis experiment. Note again that these fluxes were in opposite directions inside the membrane. This analysis can likely be further improved when we include the activity of ions outside the membrane [16].
@@
It is interesting to estimate the velocities of the water and the ions in the membrane. We assume a porosity of 34% and we consider velocities straight across the membrane. At the start of the experiment the concentrations on the two sides are 4 mM and 800 mM. The water velocity is $\sim\!300$ nm/s. The counterion velocity is quite constant across the membrane at $\sim 15$ nm/s. This is a velocity in the direction opposite to the water flow and about 5% of the water velocity. Thus the counterions have a velocity relative to the water to not be dragged along, and then a 5% extra velocity to make them move in the reference frame of the membrane. The coions have a concentration that on the high-salinity side is about a factor of 100 less than the counterions, and many more orders smaller on the low-salinity side. If velocity is calculated as a flux divided by a concentration, one finds that the velocity of the coions is much higher than of counterions, and rapidly increases towards the low-salinity side. But if coions are predominantly transported by diffusion, this interpretation of velocity must not be taken too literally – it is not the case that an individual coion accelerates while moving across the membrane [64].
@@
Figure 7: Experimental and theoretical results of salt flux, $J\mathrm{{}_{salt}}$, and water velocity, $v\mathrm{{}_{w}}$, across a cation-exchange membrane between two solutions with different salt concentration (initially $c_{\mathrm{{}_{d}}}\!=\!4$ mM and $c_{\mathrm{{}_{c}}}\!=\!800$ mM). Water flows to the high-salinity side, and salt in the other direction. There is no external pressure difference and no current. See main text for parameter settings.

### v6

@@ 5.2 Flow of solutes through membranes according to solution-friction theory
To describe the flow of ions through a porous tortuous medium, first of all we set up an expression for the chemical potential of an ion. In this section the porous medium, or material, is a membrane, so these terms are used interchangeably; we also use the term matrix for the solid phase of the membrane. We can consider multiple contributions to the chemical potential of an ion (unit J/mol) in a porous medium, adding up to [24]
@@
where the reference value $\overline{\mu}_{\text{ref},i}$ is of importance when ions take part in chemical reactions, while $RT\,\ln\left(c_{i}/c\mathrm{{}_{ref}}\right)$ relates to diffusion (ion entropy), with $c\mathrm{{}_{ref}}$ a reference concentration, for instance, $c\mathrm{{}_{ref}}=1$ mM. Furthermore, $z_{i}FV$ relates to electrostatic effects, $\overline{\mu}_{\text{aff},i}$ to a chemical affinity of the ions with the material, and $\overline{\mu}_{\text{exc},i}$ describes volumetric interactions between ions, and between ions and the porous medium [23]. When only interactions of an ion with the porous medium are of importance (and not between ions), then $\overline{\mu}_{\text{exc},i}$ is a constant factor that can be combined with $\overline{\mu}_{\text{aff},i}$.
We set up Eq. (49) for an ion inside and outside the membrane, equate the two expressions, and for a neutral solute then arrive for the distribution of a solute across the solution/membrane interface at
$\frac{c_{\text{m},i}}{c_{\infty,i}}=\exp\left(-\tfrac{1}{RT}\,\Delta\left(\overline{\mu}_{\text{aff},i}+\overline{\mu}_{\text{exc},i}\right)\right)=\Phi_{i}$ (50)
where the partition coefficient, $\Phi_{i}$, is a constant factor if the chemical affinity and excess terms do not depend (too much) on concentrations in solution and in the membrane, and where $\Delta$ refers to a difference between in the membrane and in solution [17]. The excess contribution is often described as function of an ion size-pore size ratio, $\lambda$, described for a perfectly cylindrical pore and spherical ions by $\Phi_{i}=\left(1-\lambda_{i}\right)^{2}$, but more accurate and realistic expressions are also available [14]. For ions we must extend Eq. (50) to account for their charge, and we then end up with the extended Boltzmann equation, Eq. (34).
@@
where we consider a single coordinate, x. An extension to multiple coordinates or dimensions is straightforward. We can leave out here a gradient-term related to $\overline{\mu}_{\text{aff},i}$ because this ion-membrane affinity is invariant across the membrane. And for the same reason we also leave out the excess-term that relates to volume effects (unless membrane compaction is a function of position). Thus we only have to consider a force related to diffusion, and to the electric field, i.e., electromigration.
@@
$\mathcal{F}_{\text{friction},i}=-\sum_{j}f_{i\text{-}j}^{*}\left(v_{i}-v_{j}\right)=-f_{i\text{-w}}^{*}\left(v_{i}-J\mathrm{{}_{w}}\right)-f_{i\text{-m}}^{*}\;v_{i}$ (53)
where water velocity is $J\mathrm{{}_{w}}$ and we included in the last part that the velocity of the membrane is zero, $v\mathrm{{}_{m}}\!=\!0$. We present here velocities as superficial velocities, per unit total membrane area. A more detailed analysis is the two-fluid model (TFM) which starts at interstitial velocities, and includes pore tortuosity, and also includes how ions take up space that is is not available for water  [14].
We can write the ion-water frictional coefficient as the inverse of a diffusion coefficient of the ion in the membrane, $D_{\text{m},i}$, which is the Einstein equation, $f_{i\text{-w}}^{*}=RT/{D_{\text{m},i}}$ (p. 128 in ref. [42]). This diffusion coefficient is lower than in free solution because it includes porosity and tortuosity of the membrane pores. We now obtain an extended Nernst-Planck equation for the molar flux of ions inside a membrane (Eq. (7) in ref. [5]) where convection, diffusion, and electromigration, all contribute [52, 37]
$J_{i}=K_{\text{f},i}c_{\text{m},i}J\mathrm{{}_{w}}-K_{\text{f},i}D_{\text{m},i}\left(\frac{\partial c_{\text{m},i}}{\partial x}+z_{i}c_{\text{m},i}\frac{\partial\phi}{\partial x}\right)$ (54)
and where we also incorporated that $J_{i}=c_{\text{m},i}v_{i}$. Concentrations here are those inside the membrane pores. The friction factor, or hindrance function, $K_{\text{f},i}$, is given by $K_{\text{f},i}=\left(1+f_{i\text{-m}}^{*}/f_{i\text{-w}}^{*}\right)^{-1}$, which has a value between 0 and 1, and describes that ions not only have friction with the water that directly envelopes them, but also with the membrane matrix. Eq. (54) is a general expression for ion transport in porous media, extended compared to the standard Nernst-Planck equation with convection and with ion-membrane friction, which leads to the factor $K_{\text{f},i}$.
@@ 5.3 Flow of water through membranes according to solution-friction theory
We continue with a description of transport of water through the pores of a porous medium. As discussed before, in SF theory the pores are filled with water completely, and ions have no volume, i.e., in the theory they are point charges. ref. [14]. We set up a balance of forces acting on a volume element of water in the membrane pores, with the driving force a gradient in total pressure, i.e., $-\partial P^{\text{tot}}/\partial x$ is the force acting on a volume element of water. The water has a velocity $v\mathrm{{}_{w}}$ and has friction with the membrane structure, $-f_{\text{w-m}}RT\left(v\mathrm{{}_{w}}-v\mathrm{{}_{m}}\right)$, where $v\mathrm{{}_{m}}\!=\!0$ because the membrane is immobile (ref. [42], p. 126), and water has friction with all solutes, and this friction is proportional to the concentration of solutes, which leads to $-\sum_{i}^{\text{ }}f_{i\text{-w}}^{*}c_{\text{m},i}\left(v\mathrm{{}_{w}}-v_{i}\right)$. The summation of all these forces is set to zero, resulting in [37]
$-\frac{\partial P^{\text{tot}}}{\partial x}=f_{\text{w-m}}RTv\mathrm{{}_{w}}+\sum_{i}f_{i\text{-w}}^{*}c_{\text{m},i}\left(v\mathrm{{}_{w}}-v_{i}\right)\,.$ (55)
@@
We can solve Eq. (55) at each position in the membrane, together with the extended Nernst-Planck equation, Eq. (54). A significant simplification is arrived at if we assume there is no ion-membrane friction, and thus $K_{\text{f},i}\!=\!1$ in the NP equation, Eq. (54), for all ions. Then inserting the NP equation in Eq. (55) results in
$-\frac{1}{RT}\left(\frac{\partial P^{\text{h}}}{\partial x}-\frac{\partial\Pi}{\partial x}\right)=f_{\text{w-m}}J\mathrm{{}_{w}}+\sum_{i}\left(\frac{\partial c_{\text{m},i}}{\partial x}+z_{i}c_{\text{m},i}\frac{\partial\phi}{\partial x}\right)$ (56)
@@
where we implemented that the total pressure, $P^{\text{tot}}$, is given by $P^{\text{tot}}=P^{\text{h}}-\Pi$, and $D_{\text{m},i}=RT/f_{i\text{-w}}^{*}$. Because in the ideal case that we discuss here the osmotic pressure is the total ion concentration times RT, thus $\Pi=RT\cdot\sum_{i}c_{\text{m},i}$, Eq. (56) simplifies to
$-\frac{1}{RT}\frac{\partial P^{\text{h}}}{\partial x}=f_{\text{w-m}}J\mathrm{{}_{w}}-X\,\frac{\partial\phi}{\partial x}$ (57)
@@
and thus we have an exact cancellation of the osmotic pressure acting on the water, and the diffusional forces acting on the ions, which transfer to the water via ion-water friction. We changed notation here for the water velocity, from $v\mathrm{{}_{w}}$ to $J\mathrm{{}_{w}}$. But this only occurs when the ions have no friction with other phases other than the water. In the derivation of Eq. (57) we also included local electroneutrality in the membrane, $\sum_{i}z_{i}c_{\text{m},i}+X=0$. We can rewrite Eq. (57) to [31, 50]
$J\mathrm{{}_{w}}=-k_{\text{w-m}}\left(\frac{1}{RT}\frac{\partial P^{\text{h}}}{\partial\overline{x}}-X\,\frac{\partial\phi}{\partial\overline{x}}\right)$ (58)
and thus water flow is because of a hydrostatic pressure gradient, water-membrane friction, and an electrostatic body force term, which entered this force balance via water-ion interaction. For an uncharged membrane, $X\!=\!0$, and then Eq. (58) simplifies to the Darcy equation, Eq. (26). Thus water flows across an uncharged membrane not by diffusion but because of an internal pressure gradient [35, 60, 61, 27]. For a constant membrane charge, it can be easily integrated between the two sides of the membrane. Let it be reminded that the last three equations all assumed $K_{\text{f},i}\!=\!1$.
@@
The final element of transport theory through a porous structure filled with a continuum fluid such as water, with dissolved solutes, is mechanical equilibrium at the interface between the porous medium and the outside solution. This was already discussed in section 3. Here we derive the mechanical equilibrium condition on the basis on Eq. (56), because with all gradient terms much larger in a thin interfacial region than the terms proportional to velocities (or in other words, the structure of the interfacial region is not significantly modified by flows of ions and water across it), we obtain the result that the total pressure $P^{\text{tot}}=P^{\text{h}}-\Pi$ is invariant across any thin interface [36, 50, 46], and this leads to the earlier conclusion that the change in osmotic pressure $\Pi$ (either up or down) equals the change in hydrostatic pressure $P^{\text{h}}$ across an interface.
@@ 5.4 Counterfluxes of salt and water in osmosis
To illustrate SF theory, and to show how well it works, we discuss results of an accurate fit of SF theory to data of the flow of water and salt across an ion-exchange membrane. This experiment is called osmosis, and water and salt flow because of a difference in salt concentration between the two outside solutions on either side of the membrane, without pressure differences or current. Results of this experiment are presented in Fig. 6 as function of the concentration on the low-salinity side, $c\mathrm{{}_{d}}$, which changes in time from 4 mM to 130 mM. During that time the salt concentration on the concentrate side decreases from 800 to 525 mM. Over the 35 hr of the experiment, the volume on the diluate side drops by 40% because water flows to the concentrated side [14]. Thus, in this experiment water flows to the high-salinity side, while salt flows in the other direction. Consequently, inside the membrane these two flows are in opposite direction. This makes this an experiment that is not easy to reproduce accurately in a theoretical model, i.e., it is a very good test of a mass transport theory.
@@
When we apply SF theory, with cation and anion fluxes the same (which is because the electrical current through the membrane is zero), and with equal diffusion coefficients of the ions in the membrane, we find that it works excellently, and results in the very good fit to the data shown in Fig. 6. The salt flux, $J\mathrm{{}_{salt}}$, follows from solving Eq. (54) while implementing that the flux of cations is equal to that of anions, and then we integrate. This leads to Eq. (5) in ref. [3] which we solve assuming $c\mathrm{{}_{T,m}}$ changes linearly, after which we implement Eq. (46). For the partition coefficient, we use $\Phi_{i}\!=\!0.82$ [63]. The water flux is calculated from Eq. (55) which after integration results in Eq. (S17) in ref. [3]. Membrane charge density is $X\!=\!-5.1$ M. To fit the model to the data, we derive $k\mathrm{{}_{m}}\!=\!2.25$ LMH (for a membrane thickness of $L\mathrm{{}_{m}}=120$ $\mu$m, this corresponds to a diffusion coefficient in the membrane about 20$\times$ less than in free solution), $K_{\text{f},i}\!=\!0.50$, and for the water permeability we use $A\!=\!9.2$ mLMH/bar. Values for $k\mathrm{{}_{m}}$ and $A$ are in line with earlier reported values for ED membranes [53]. Thus, we can conclude that SF theory, which includes chemical and mechanical equilibrium on solution/membrane interfaces, is very well capable of describing fluxes of water and salt in the osmosis experiment. Note again that these fluxes were in opposite directions inside the membrane. This analysis can likely be further improved when we include the activity of ions outside the membrane. Inside the membrane, they are rather constant, thus in effect included in $\Phi_{i}$ [17].
@@
It is interesting to estimate the velocities of water and ions in the membrane. We assume a porosity of 34% and consider velocities straight across the membrane. At the start of the experiment the concentrations on the two sides are 4 mM and 800 mM. The water velocity is $\sim\!300$ nm/s. The counterion velocity is quite constant across the membrane at $\sim 15$ nm/s. This is a velocity in the direction opposite to the water flow, about 5% of the water velocity. Thus the counterions have a velocity relative to the water to not be dragged along, and then a 5% extra velocity to make them move in the reference frame of the membrane. The coions have a concentration that on the high-salinity side is about a factor of 100 less than the counterions, and many more orders smaller on the low-salinity side. If velocity is calculated as a flux divided by a concentration, one finds that the velocity of the coions is much higher than of counterions, and rapidly increases towards the low-salinity side. But if coions are predominantly transported by diffusion, this interpretation of velocity must not be taken too literally – it is not the case that an individual coion accelerates while moving across the membrane [64].
@@
Figure 6: Experimental and theoretical results of salt flux, $J\mathrm{{}_{salt}}$, and water velocity, $J\mathrm{{}_{w}}$, across a cation-exchange membrane between two solutions with different salt concentration (initially $c_{\mathrm{{}_{d}}}\!=\!4$ mM and $c_{\mathrm{{}_{c}}}\!=\!800$ mM). Water flows to the high-salinity side, and salt in the other direction. There is no external pressure difference and no current. See main text for parameter settings.